# A Damage-Driven Model for Duchenne Muscular Dystrophy: Early-Stage Dynamics and Invasion Thresholds

### 31 Mar 2026 v2

@@ 4.1.4 Boundedness of $\Phi$
Integrating the equations for $h$ and $d$ in (4.30) and summing, we obtain
@@
Here $|\Omega|$ denotes the measure of the domain. Applying Gronwall’s lemma and using that $h_{in}+d_{in}\leq 1$, we conclude that $h$ and $d$ are bounded in $L^{1}(\Omega)$. Integrating the equations for $m$ and $c$ in (4.30), we obtain
@@
Proof. Estimate (4.35) follows from the equation for $c$ in (4.30), using maximal regularity for the heat equation together with (4.31). Estimate (4.36) then follows from (4.35) by the embedding Lemma 6 (see [5]).
@@
Let $(h,d,m,c)\in\Phi$. Under the hypotheses of Theorem 5, for any $1\leq p<\infty$, there exists a constant $H_{p}$, depending on $p$, $\Omega$, $T$, $n$, the initial data and the parameters of the system, but not on $\lambda$, such that
@@
[…]
The function $\psi_{D}$ is non-negative. Moreover, by maximal regularity for the heat semigroup [5, 21] and Lemma 6, we have
@@
where $q_{1}=(n+2)p{{}^{\prime}}/(n+2-2p{{}^{\prime}})$, $q_{2}=(n+2)p{{}^{\prime}}/(n+2-p{{}^{\prime}})$, and $C_{p{{}^{\prime}}}$ is a positive constant depending on $\Omega$, $T$, $n$, and $p{{}^{\prime}}$. Multiplying the equation by $\theta$ and integrating over space and time, and using (4.38) with $D=D_{m}$ together with integration by parts and (4.30), we obtain
@@
Analogously, multiplying $\theta$ by $h+d$ and integrating over space and time, and using (4.38) with $D=1$ and $D=D_{d}$ together with integration by parts and (4.30), we obtain
@@
We begin with the initial term in (4.40). Using Hölder’s inequality together with (4.39) (see [5]), we obtain
@@
The same argument applied to the first term in (4.41) yields
@@
We now estimate the last term in (4.41). Using (4.39), we obtain
@@
We next estimate the last term in (4.40). Following [5, 12], we introduce $\vartheta=1-\frac{2}{n+2}$, so that $0<\vartheta<1$. Using Hölder’s inequality together with (4.31), (4.39) and Lemma 6, we obtain
@@
The same estimate applied to (4.41) yields
@@
It remains to estimate the remaining terms in (4.40) and (4.41).
We fix $s$ and $\xi$ in $(0,1)$ such that
@@
Using (4.47) and (4.48), together with Hölder’s and Young’s inequalities, and the estimates (4.31) and (4.36), we obtain
@@
Collecting the estimates (4.42), (4.45), and (4.49) in (4.40), together with (4.43), (4.46) and (4.50) in (4.41), we obtain
@@
Since the above inequality holds for every $\theta\geq 0$ with $\|\theta\|_{L^{p{{}^{\prime}}}(\Omega_{T})}\leq 1$, by duality in $L^{p}(\Omega_{T})$ we obtain
$\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}\leq C_{1}+C_{2}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\vartheta}+C_{3}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{s}+C_{4}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\xi}\,$,
@@
Under the hypotheses of Theorem 5, the set $\Phi$ is bounded in $C^{0,1}]^{4}$. In particular, if $(h,d,m,c)\in\Phi$, then
@@
Proof. Since $p>n+2$, the space $W^{1,2}_{p}(\Omega_{T})$ is compactly embedded in $C^{0,1}$ (see Lemma 6). Applying maximal regularity to the equation for $c$ in (4.30) and using (4.37), we obtain
@@
Repeating the same argument for the $m$-equation in (4.30), we obtain
@@
Proof. Suppose that there exist $t\in[0,T]$ and $x_{0}\in\overline{\Omega}$ such that $h(t,x_{0})+d(t,x_{0})>1$. Let $A$ be a neighborhood of $x_{0}$ such that $h(t,x)+d(t,x)>1$ for all $x\in A\subset\overline{\Omega}$. Then, for every ball $B_{r}$ of radius $r$, centered at $x_{0}$ and contained in $A$, we have
$\frac{1}{|B_{r}|}\int_{B_{r}}(f+g)\,dx>1\,.$ (4.57)
On the other hand, using the equations for $h$ and $d$ in (4.30) and the condition $h_{in}+d_{in}\leq 1$, we obtain
$\partial_{t}\int_{B_{r}}(h+d)dx\leq t\int_{\partial B_{r}}\left(\nabla h\right)\cdot\mathbf{n}\,ds+t\,D_{d}\int_{\partial B_{r}}\left(\nabla d\right)\cdot\mathbf{n}\,ds.$ (4.58)
This inequality holds for every ball $B_{r}$ of radius $r$, centered at $x_{0}$ and contained in $A$.
From (4.58), and using that $h,d\in C^{0,1}$, we deduce that
$\frac{1}{|B_{r}|}\int_{B_{r}}(h+d)\,dx\leq 1+o(1)\quad\text{as }r\to 0,$
which contradicts (4.57). This concludes the proof.
@@
[…]
@@ 4.1.5 Existence of the solution
Moreover, there exists a constant $C$, depending on $\Omega$, $T$, $n$, the parameters of the system and $\|(h_{in},d_{in},m_{in},c_{in})\|_{[W^{2-2/\bar{p},\bar{p}}(\Omega)]^{4}}$, such that
@@ 4.2 Uniqueness
$\frac{d}{dt}\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}+2D_{c}\|\nabla(c_{1}-c_{2})\|^{2}_{L^{2}(\Omega)}\leq r\|d_{1}-d_{2}\|^{2}_{L^{2}(\Omega)}\,$, (4.61)
$\frac{d}{dt}\|m_{1}-m_{2}\|^{2}_{L^{2}(\Omega)}+2D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}\leq\nu\|d_{1}-d_{2}\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{8D_{m}}\|m_{1}\|^{2}_{C^{0,1}}\|\nabla(c_{1}-c_{2})\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{8D_{m}}\|c_{2}\|^{2}_{C^{0,1}}\|m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\kappa^{2}\chi_{0}^{2}}{8D_{m}}\|m_{2}\|^{2}_{C^{0,1}}\|c_{2}\|^{2}_{C^{0,1}}\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}\,$, (4.62)
@@
Denoting by $A=\frac{3}{16}\frac{\chi_{0}^{2}}{D_{m}\,D_{d}}\|m_{1}\|^{2}_{C^{0,1}}$, we complete the proof applying the Gronwall’s lemma to
@@ 4.3 Uniform-in-time estimates
We prove three preliminary Lemmas where we establish, respectively, uniform in time bounds in $L^{1}(\Omega)$ for the solution, uniform in time bounds for the gradient of $c$ and finally we state uniform in time bounds in $L^{p}(\Omega)$ for the solution. We conclude this sub Section with Proposition 20 where we prove the uniform in time estimate of the solution in $W^{1,\infty}(\Omega)$ (see [5]).
@@
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}_{1}$, depending on $\Omega$, $n$ the parameters of the system and the initial data, such that
@@
and applying Gronwall’s lemma, we obtain (4.64) and (4.65) for $h$ and $d$.
Integrating the equation of $m$ and $c$ in (4.1), we have:
@@
$\frac{d}{dt}\int_{\Omega}c\,dx\leq-\mu\int_{\Omega}c\,dx+r\|d\|_{L^{1}(\Omega)}\,$,
and (4.64) and (4.65) for $m$ and $4c$ are proved applying Gronwall’s lemma.
To prove the next two lemmas, following [5], we define a smooth positive cutoff function $\upsilon\in C^{\infty}\left(\mathbb{R};[0,1]\right)$ such that $\sup_{s\in\mathbb{R}}|\upsilon(s)|\leq 1$ and $\upsilon(s)=1$ for $s\geq 1$ and $\upsilon(s)=0$ for $s\leq 0$ and $|\upsilon{{}^{\prime}}(s)|\leq 2$ for all $s\in\mathbb{R}$. For any $\tau\in\mathbb{N}$, we define $\upsilon_{\tau}(\cdot)=\upsilon(\cdot-\tau)$. The idea is to multiply the equations (2.3) by $\upsilon_{\tau}$ and repeating the steps using in Lemmas 11 and 12 to obtain new estimates in $\Omega_{\tau,\tau+1}$ which are independent of $\tau$.
@@
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}^{\ast}$, depending on $\Omega$, $n$ the parameters of the system and the initial data, but not on $\tau$, such that
@@
Multiplying the $c$-equation by $\upsilon_{\tau}$, considering that $\upsilon_{\tau}c=0$ in $(x,\tau)$ and using the regularity property for the heat equation, we have
$\|\upsilon_{\tau}\nabla c\|_{L^{\frac{n+2}{n+1}}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}{{}^{\prime}}c+\upsilon_{\tau}c-\mu\upsilon_{\tau}d\|_{L^{1}(\Omega_{\tau,\tau+2})}\,$, (4.69)
where $\overline{C}^{\ast}$ depends on $\Omega$, $n$ and the parameters of the system, but not on $\tau$. Estimates (4.67) and (4.68) follows using (4.64) and (4.65) and the properties of $\upsilon_{\tau}$.
@@
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}_{p}$, depending on $\Omega$, $p$, $n$, the parameters of the system and the initial data, but not on $\tau$, such that
@@
Multiply $\theta$ by $\upsilon_{\tau}m$ and integrate in space and time, using equation (4.38) in $\Omega_{\tau,\tau+2}$, integrating by parts and using (4.1), we obtain:
@@
Analogously, multiply $\theta$ by $\upsilon_{\tau}(h+d)$ integrate in space and time, using equations (4.38) in $\Omega_{\tau,\tau+2}$, integrating by parts and using (4.1), we obtain:
@@
We can estimate each terms in (4.72), (4.73) and (4.74) following the proof in Lemma 12 and using (4.64) and (4.65) and (4.67) and (4.68), obtaining
@@
Under the hypotheses of Theorem 5, let $(h,d,m,c)$ be the unique solution to system (4.1) with Neumann boundary conditions (4.3) and initial data (4.4), constructed in Propositions 15 and 16. Then there exists a constant $\overline{C}>0$, depending on $\Omega$, $h_{in}$, $d_{in}$, $m_{in}$, $c_{in}$ and on the parameters of the system, such that the solution satisfies:
@@
Proof. Using the embedding properties stated in Lemma 6 together with the maximal regularity theory for the heat equation, and taking $p>n+2$, we obtain
$\|\upsilon_{\tau}c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}c\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}{{}^{\prime}}c-\upsilon_{\tau}\mu c+r\upsilon_{\tau}\,d\|_{L^{p}(\Omega_{\tau,\tau+2})}\,$, (4.79)
where $\overline{C}^{\ast}$ depends on $\Omega$, $n$, and the parameters of the system, but not on $\tau$. Using (4.70) and (4.71) and the definition of $\upsilon_{\tau}$, the Lemma is proved for $c$.
Analogously for $h$ and $d$, as
$\|\upsilon_{\tau}h\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla h\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}h\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|(\upsilon_{\tau}{{}^{\prime}}-\upsilon_{\tau})h\|_{L^{p}(\Omega_{\tau,\tau+2})}\\ +\overline{C}^{\ast}\|\upsilon_{\tau}(\sigma+\rho\,d)(1-h-d)-\upsilon_{\tau}\,A(m,c)h\|_{L^{p}(\Omega_{\tau,\tau+2})}\,$, (4.80)
@@
$\|\upsilon_{\tau}d\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla d\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}d\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|(\upsilon_{\tau}{{}^{\prime}}-\delta\upsilon_{\tau})d\|_{L^{p}(\Omega_{\tau,\tau+2})}\\ +\overline{C}^{\ast}\|\upsilon_{\tau}\,A(m,c)h-\upsilon_{\tau}\,\rho\,d(1-h-d)\|_{L^{p}(\Omega_{\tau,\tau+2})}\,.$ (4.81)
For the $m$-equation, following [5], we write
$\upsilon_{\tau}m=\int_{\tau}^{t}e^{-D_{m}\Delta(t-\tau-s)}\left(\upsilon_{\tau}\nu(d-m)-\upsilon_{\tau}{{}^{\prime}}m-\nabla\cdot\left(\chi(m,c)\nabla c\right)\right)ds\,$, (4.82)

### 03 Apr 2026 v3

@@ 4.1.4 Boundedness of $\Phi$
Integrating the equations for $h$ and $d$ in (4.30) and summing, we obtain:
@@
Here $|\Omega|$ denotes the measure of the domain. Applying Gronwall’s lemma and using that $h_{in}+d_{in}\leq 1$, we conclude that $h$ and $d$ are bounded in $L^{1}(\Omega)$. Integrating the equations for $m$ and $c$ in (4.30), we obtain:
@@
Proof. From the equation for $c$ in (4.30), we may write
$\nabla c(t)=e^{-\mu t}\nabla e^{tD_{c}\Delta}(\lambda c_{in})+r\lambda\int_{0}^{t}e^{-\mu(t-s)}\nabla e^{(t-s)D_{c}\Delta}d(s)\,ds.$
Using classical $L^{p}$-$L^{q}$ estimates (see [14]), we obtain
$\|\nabla c(t)\|_{L^{\frac{n+1}{n+2}}(\Omega)}\leq e^{-\mu t}\|\nabla e^{tD_{c}\Delta}c_{in}\|_{L^{\frac{n+1}{n+2}}(\Omega)}+Cr\int_{0}^{t}e^{-\mu(t-s)}(t-s)^{-\frac{n+1}{n+2}}\|d(s)\|_{L^{1}(\Omega)}\,ds.$
The first term is bounded uniformly on $[0,T]$ by the regularity of the initial datum, while the second term by (4.31). Therefore,
$\sup_{t\in[0,T]}\|\nabla c(t)\|_{L^{\frac{n+1}{n+2}}(\Omega)}\leq C,$
with $C$ independent of $\lambda$ and this concludes the proof.
@@
Let $(h,d,m,c)\in\Phi$. Under the hypotheses of Theorem 5, for any $1\leq p<\infty$, there exists a constant $H_{p}$, depending on $p$, $\Omega$, $T$, $n$, the initial data and the parameters of the system, but not on $\lambda$, such that:
@@
$\begin{cases}\begin{array}[]{l}\frac{\partial}{\partial t}\psi_{D}+D\Delta\psi_{D}=-\theta\qquad(t,x)\in\Omega_{T}\\ \nabla\psi_{D}\cdot\mathbf{n}=0\qquad\qquad(t,s)\in(0,T)\times\partial\Omega\\ \psi_{D}(T,x)=0\qquad\qquad x\in\Omega\end{array}\end{cases}$ (4.37)
The function $\psi_{D}$ is non-negative. Moreover, by maximal regularity for the heat semigroup [6, 22] and Lemma 6, we have:
@@
where $q_{1}=(n+2)p{{}^{\prime}}/(n+2-2p{{}^{\prime}})$, $q_{2}=(n+2)p{{}^{\prime}}/(n+2-p{{}^{\prime}})$ and $C_{p{{}^{\prime}}}$ is a positive constant depending on $\Omega$, $T$, $n$ and $p{{}^{\prime}}$. Multiplying the equation by $\theta$, integrating over space and time and applying (4.37) with $D=D_{m}$ together with integration by parts and (4.30), yields:
@@
Proceeding as above, multiplying $\theta$ by $h+d$, integrating over space and time, and applying (4.37) with $D=1$ and $D=D_{d}$ together with integration by parts and (4.30), yields:
@@
We begin with the initial term in (4.39). Using Hölder’s inequality together with (4.38) (see [6]), we obtain:
@@
The same argument applied to the first term in (4.40) gives:
@@
We now estimate the last term in (4.40). Using (4.38), we have:
@@
We next estimate the last term in (4.39). Following [6, 13], we introduce $\vartheta=1-\frac{2}{n+2}$, so that $0<\vartheta<1$. Using Hölder’s inequality together with (4.31), (4.38) and Lemma 6, we obtain:
@@
The same estimate applied to (4.40) yields:
@@
It remains to estimate the terms in (4.39) and (4.40).
We fix $s$ and $\xi$ in $(0,1)$ such that:
@@
Using (4.46) and (4.47), together with Hölder’s and Young’s inequalities and the estimates (4.31) and (4.35), we have:
@@
Collecting the estimates (4.41), (4.44) and (4.48) in (4.39), together with (4.42), (4.45) and (4.49) in (4.40), we obtain:
@@
Since the above inequality holds for every $\theta\geq 0$ with $\|\theta\|_{L^{p{{}^{\prime}}}(\Omega_{T})}\leq 1$, duality in $L^{p}(\Omega_{T})$ yields:
$\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}$ $\leq C_{1}+C_{2}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\vartheta}+C_{3}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{s}$
$+C_{4}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\xi}\,$,
@@
Under the hypotheses of Theorem 5, the set $\Phi$ is bounded in $C^{0,1}]^{4}$. In particular, if $(h,d,m,c)\in\Phi$, then:
@@
Proof. Since $p>n+2$, the space $W^{1,2}_{p}(\Omega_{T})$ is compactly embedded in $C^{0,1}$ (see Lemma 6). Applying maximal regularity to the equation for $c$ in (4.30) and using (4.36), we obtain:
@@
Repeating the same argument for the $m$-equation in (4.30), we have:
@@
Proof. From (4.30), the $2\times 2$ parabolic semilinear system for $h$ and $d$ has a diagonal diffusion. Arguing as in Theorem 1 and considering Lemma 9, the reaction field is inward pointing on the boundary face $h+d=1$, and on $h=0$ and $d=0$. Applying the invariant-region theorem for semilinear parabolic systems (see [4, 34]) yields that the convex bounded domain
$\Sigma=\{(h,d)\in\mathbb{R}^{2}:\ h\geq 0,\ d\geq 0,\ h+d\leq 1\}$
is positively invariant and the Lemma is proved.
@@
$\begin{cases}\displaystyle\partial_{t}h=\Delta h-h-A(m,c)h+\lambda\sigma(1-(h+d))+\rho d(1-(h+d)),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}d=D_{d}\Delta d-(\delta+\rho)d+A(m,c)h+\rho\lambda d(h+d),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m,c)\nabla c\right)+\nu(\lambda d-m),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}c=D_{c}\Delta c-\mu c+r\lambda d\,,&(t,x)\in\Omega_{T}\end{cases}$ (4.56)
@@ 4.1.5 Existence of the solution
Moreover, there exists a constant $C$, depending on $\Omega$, $T$, $n$, the parameters of the system and $\|(h_{in},d_{in},m_{in},c_{in})\|_{[W^{2-2/\bar{p},\bar{p}}(\Omega)]^{4}}$, such that:
@@ 4.2 Uniqueness
$\frac{d}{dt}\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}+2D_{c}\|\nabla(c_{1}-c_{2})\|^{2}_{L^{2}(\Omega)}\leq r\|d_{1}-d_{2}\|^{2}_{L^{2}(\Omega)}+r\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}-2\mu\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}\,$, (4.58)
$\frac{d}{dt}\|m_{1}-m_{2}\|^{2}_{L^{2}(\Omega)}+2D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}\leq\nu\|d_{1}-d_{2}\|^{2}_{L^{2}(\Omega)}-\nu\|m_{1}-m_{2}\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{2D_{m}}\|m_{1}\|^{2}_{C^{0,1}}\|\nabla(c_{1}-c_{2})\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{2D_{m}}\|c_{2}\|^{2}_{C^{0,1}}\|m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{2\kappa^{2}D_{m}}\|m_{2}\|^{2}_{C^{0,1}}\|c_{2}\|^{2}_{C^{0,1}}\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}\,$, (4.59)
@@
Denoting by $A=\frac{3}{4}\frac{\chi_{0}^{2}}{D_{m}\,D_{c}}\|m_{1}\|^{2}_{C^{0,1}}$, we complete the proof applying the Gronwall’s lemma to:
@@ 4.3 Uniform-in-time estimates
We prove three preliminary Lemmas where we establish, respectively, uniform in time bounds in $L^{1}(\Omega)$ for the solution, uniform in time bounds for the gradient of $c$ and finally we state uniform in time bounds in $L^{p}(\Omega)$ for the solution. We conclude this section with Proposition 20, where we prove the uniform in time estimate of the solution in $W^{1,\infty}(\Omega)$ (see [6]).
@@
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}_{1}$, depending on $\Omega$, $n$, the parameters of the system and the initial data, such that:
@@
Applying Gronwall’s lemma, we have (4.61) and (4.62) for $h$ and $d$.
Integrating the equation of $m$ and $c$ in (4.1), we obtain:
@@
$\frac{d}{dt}\int_{\Omega}c\,dx\leq-\mu\int_{\Omega}c\,dx+r\|d\|_{L^{1}(\Omega)}\,.$
Moreover, (4.61) and (4.62) for $m$ and $c$ are proved applying Gronwall’s lemma.
To prove the next two lemmas, following [6], we define a smooth positive cutoff function $\upsilon\in C^{\infty}\left(\mathbb{R};[0,1]\right)$ such that $\sup_{s\in\mathbb{R}}|\upsilon(s)|\leq 1$ and $\upsilon(s)=1$ for $s\geq 1$ and $\upsilon(s)=0$ for $s\leq 0$ and $|\upsilon{{}^{\prime}}(s)|\leq 2$ for all $s\in\mathbb{R}$. For any $\tau\in\mathbb{N}$, we define $\upsilon_{\tau}(\cdot)=\upsilon(\cdot-\tau)$. The idea is to multiply the equations (4.1) by $\upsilon_{\tau}$ and repeating the steps used in Lemmas 11 and 12 to obtain new estimates in $\Omega_{\tau,\tau+1}$ which are independent of $\tau$.
@@
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}^{\ast}$, depending on $\Omega$, $n$, the parameters of the system and the initial data, but not on $\tau$, such that:
@@
Repeating the same arguments in Lemma 11 for the equation of $\upsilon_{\tau}\,c$, considering that $\upsilon_{\tau}c=0$ in $(x,\tau)$ and using (4.61) and (4.62) and the properties of $\upsilon_{\tau}$, we obtain (4.64) and (4.65).
@@
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}_{p}$, depending on $\Omega$, $p$, $n$, the parameters of the system and the initial data, but not on $\tau$, such that:
@@
Multiplying $\theta$ by $\upsilon_{\tau}m$, integrating in space and time and using equation (4.37) in $\Omega_{\tau,\tau+2}$, integrating thus by parts and using (4.1), we obtain:
@@
Proceeding as above, we multiply $\theta$ by $\upsilon_{\tau}(h+d)$, integrate over space and time, and apply (4.37) in $\Omega_{\tau,\tau+2}$ together with integration by parts and (4.1), which yields:
@@
We can estimate each terms in (4.68), (4.69) and (4.70) following the proof in Lemma 12 and using (4.61), (4.62), (4.64) and (4.65), obtaining:
@@
Under the hypotheses of Theorem 5, let $(h,d,m,c)$ be the unique solution to system (4.1) with Neumann boundary conditions (4.3) and initial data (4.4), constructed in Propositions 15 and 16. Then, there exists a constant $\overline{C}>0$, depending on $\Omega$, $h_{in}$, $d_{in}$, $m_{in}$, $c_{in}$ and on the parameters of the system, such that the solution satisfies:
@@
Proof. Using the embedding properties stated in Lemma 6 together with the maximal regularity theory for the heat equation and taking $p>n+2$, we obtain:
$\|\upsilon_{\tau}c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}c\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}$
$\leq\overline{C}^{\ast}\|\upsilon_{\tau}{{}^{\prime}}c-\upsilon_{\tau}\mu c+r\upsilon_{\tau}\,d\|_{L^{p}(\Omega_{\tau,\tau+2})}\,$, (4.75)
where $\overline{C}^{\ast}$ depends on $\Omega$, $n$ and the parameters of the system, but not on $\tau$. Using (4.66), (4.67) and the definition of $\upsilon_{\tau}$, the Lemma is proved for $c$.
Analogously for $h$ and $d$, as:
$\|\upsilon_{\tau}h\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla h\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}h\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}\leq$
$\overline{C}^{\ast}\|(\upsilon_{\tau}{{}^{\prime}}-\upsilon_{\tau})h\|_{L^{p}(\Omega_{\tau,\tau+2})}+\overline{C}^{\ast}\|\upsilon_{\tau}(\sigma+\rho\,d)(1-h-d)-\upsilon_{\tau}\,A(m,c)h\|_{L^{p}(\Omega_{\tau,\tau+2})}\,$
@@
$\|\upsilon_{\tau}d\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla d\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}d\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}\leq$
$\overline{C}^{\ast}\|(\upsilon_{\tau}{{}^{\prime}}-\delta\upsilon_{\tau})d\|_{L^{p}(\Omega_{\tau,\tau+2})}+\overline{C}^{\ast}\|\upsilon_{\tau}\,A(m,c)h-\upsilon_{\tau}\,\rho\,d(1-h-d)\|_{L^{p}(\Omega_{\tau,\tau+2})}\,.$
For the $m$-equation, following [6], we write:
$\upsilon_{\tau}m=\int_{\tau}^{t}e^{-D_{m}\Delta(t-\tau-s)}\left(\upsilon_{\tau}\nu(d-m)+\upsilon_{\tau}{{}^{\prime}}m-\upsilon_{\tau}\nabla\cdot\left(\chi(m,c)\nabla c\right)\right)ds\,$, (4.76)